Calculate the visitation rates per 1,000 population in each zone. These are the total visits per year designated by each zone, divided by the zone’s population in thousands. An example is shown below. Use Microsoft Excel (or something similar) to calculate the rates.

Visitation Rates per 1,000 Population Total Visits/Year 400 400 Zone Population 1,000 2,000 4,000 8,000 Zone Visits/1,000 0 400 4007000200 100 400 50 Beyond 3 Total Visits 1,600
Answer #1

Suppose there are 5 zones, zones are named as a, b, c, d and e. Each zone has a population of P1, P2,P3, P4 and P5. and total visit from each zone is V1, V2,V3, V4 and V5.

In excel sheet, this will look like -

Zone Total Visits/Year Zone Population Visits/1000
a V1 P1 1000*V1/P1
b V2 P2 1000*V2/P2
c V3 P3 1000*V3/P3
d V4 P4 1000*V4/P4
e V5 P5 1000*V5/P5
Total V1+V2+V3+V4+V5 P1+P2+P3+P4+P5 1000*(V1+V2+V3+V4+V5)/(P1+P2+P3+P4+P5)

In the above example given in the question, the total visitation rate would be 1000*1600/(1000+2000+4000+8000)=1600/15=106.67